Kutztown University Department of Mathematics

The Second Annual Thomas L. Pirnot Lecture in Mathematics

3:30 p.m.

April 14, 2009

Lytle Hall 228

Ramsey’s Theorem:

The Extremal Limits of Party Planning

Dr. Michael Picollelli

Assistant Professor of Mathematics Department of Mathematics LaFayette College

Abstract

Out of any six people, there are three mutual acquaintances or three mutual strangers.” This simple-looking statement may appear as little more than a recreational problem for mathematics contests (where it did once appear).

In fact, it is the very tip of a mathematical iceberg known as Ramsey Theory, a field which has played a pivotal role in the development of modern combinatorics.

In this talk, we will explore the fundamental result in this area, Ramsey’s Theorem, which guarantees, among other things, that a large enough party will be very good or very bad.

Refreshments served between 3:00 and 3:30 p.m.

The talk begins at 3:30 p.m.

Kutztown Univ. of Pennsylvania Dept. of Mathematics Dr. Paul S. Ache III, Chair